OpenStudy (anonymous):

can you please help me solve 2x-3/x + 4 =x-3/x

OpenStudy (anonymous):

1) 2x-3/x + 4 =x-3/x 2) 2x-3/x + 4 - x + 3/x= 0 3) 2x + 4 - x = 0 4) 2x - x = -4 5) x = -4

OpenStudy (anonymous):

\[\frac{2x-3}{x+4} = \frac{x-3}{x}\] Is this it?

OpenStudy (anonymous):

yes thank you